solve the system using either method of subsititution or elimination 3x+5y+4z=13,5x+2y+3z=-9,6x+3y+4z=-8
Answer by mananth(16946)

3x+5y+4z=13 --------------1
5x+2y+3z=-9 --------------2
6x+3y+ 4z=-8 --------------3

consider equation 1 &2 Eliminate y
Multiply 1 by -2
Multiply 2 by 5
we get
-6x-10y-8z=-26
25x+10y+15z=-45
Add the two
19x+0y +7z=-71 -------------4
consider equation 2 & 3 Eliminate y
Multiply 2 by -3
Multiply 3 by 2
we get
-15x-6y -9z=27
12x+6y+ 8z=-8
Add the two
-3x+0y-1z=19 -------------5
Consider (4) & (5) Eliminate x
Multiply 4 by 3
Multiply (5) by 19
we get 2
57x+21z =-213
-57x-19z=361
Add the two
0x+2z=148
/ 2
z=74

Plug the value of z in 5
-3x-74= 19
-3x=93
x=-31
plug value of x & z in 1
-93+5y+ 296=13
5y=93-296+13
y= -38